OpenStudy (anonymous):

Find the equation of a line parallel to 2x+2y=7 that goes through the point (3, 1/5) .....

OpenStudy (aroub):

2y=-2x+7 y=-x+7/2 y=mx+b, m=-1 , (3,1/5) 1/5=-1(3)+b 1/5=-3+b 1/5+3/1=b 1/5+15/5=b 16/5=b y=-x+16/5